OpenStudy (anonymous):

how long would it take to get 3 miles while driving 20 miles an hour

Directrix (directrix):

I get 9 minutes but that does not "sound" right.

OpenStudy (hoblos):

it is right @Directrix

Directrix (directrix):

20 miles is to 60 minutes as 3 miles is to M minutes.

Directrix (directrix):

1/3 = 3/M M = 9 minutes